# On The Fragility of Benchmark Contamination Detection in Reasoning Models

### v1

@@ 1 Introduction
Accordingly, various benchmark contamination detection methods have been proposed to determine whether specific benchmarks were used during training (Yeom et al., 2018; Mattern et al., 2023; Shi et al., 2023; Dong et al., 2024; Tu et al., 2024), based on the assumption that contamination in LLMs primarily involves memorizing the benchmark data (Wu et al., 2025). These methods rely on separability in some distributions between members (i.e., seen samples during contamination) and non-members (i.e., unseen samples). However, as LLMs have started to evolve into Large Reasoning Models (LRMs) (Guo et al., 2025; Jaech et al., 2024), benchmark contamination detection faces two key challenges: (1) LRMs rely on chain-of-thought (CoT) reasoning to reach final answers, but model developers would not release their training CoT data, and contamination detectors typically only have access to question-answer pairs without the intermediate reasoning steps used during training. This absence of training sequences makes detection substantially more challenging. (2) LRMs primarily acquire reasoning abilities during two stages: SFT and RL. This potentially provides developers with opportunities to manipulate leaderboard performance by strategically contaminating benchmarks in the earlier stage, while evading detection methods through subsequent training. Given these challenges, the effectiveness of existing detection methods against LRM contamination remains uncertain.
@@ 2 Related Works
Benchmark Contamination Concealment. Model developers hope to conceal contamination evidence while still having performance inflation. Prior work has explored evading detection through benchmark augmentation, such as rephrasing solutions with strong LLMs (Dekoninck et al., 2024; Samuel et al., 2024), but in LRM settings, most benchmarks only have question–answer pairs without step-by-step solutions, making such methods inapplicable. (Bordt et al., 2024) explores from the training dynamic perspective, showing that performance inflation due to contamination diminishes as pre-training progresses. To our knowledge, we are the first to investigate contamination concealment at the algorithmic level.
@@ 3.1 GRPO Conceals Benchmark Contamination
Contamination Inflation Mainly Comes From SFT. We evaluate multiple contamination scenarios that may happen during SFT and RL and summarize the empirical results in Tab. 1. Results show that clean SFT training yields an 11.30% improvement in pass@1 performance, while SFT contamination further inflates results by an additional 8.82% on average across six benchmarks. In contrast, RL contamination, despite introducing the benchmark questions and giving rewards based on the model-generated responses, shows no significant difference compared to using a clean RL training set. These findings indicate that after extensive SFT training, introducing benchmark samples during the RL process provides little performance benefit.
Table 1: Pass@1 (%) under different contamination scenarios when the base model evolves into LRMs. Empirical results demonstrate that contamination inflation mainly comes from the SFT stage. “/” means not used, and “Mem” denotes members. We first train the base model with SFT and then RL. The last row denotes the results of the base model.
@@
Table 2: AUROC (%) of contamination detection approaches evaluated starting from an SFT-contaminated model w/o RL to subsequently trained with GRPO. Results demonstrate that after GRPO, AUROC decreases across all the benchmarks and detection approaches. $\Delta$ measures the difference with the SFT-contaminated model w/o RL. Higher AUROC, better detection performance. Each AUROC is averaged over detection scores from 8 rollouts.
@@
To understand whether current contamination detection methods can still successfully detect contamination in LRMs, and whether RL training can alter the signals exploited by contamination detectors, we evaluate SFT-contaminated models before and after GRPO. Tab. 2 reveals systematic shifts in AUROC across diverse detection methods. Our analysis highlights three key observations:
SFT contamination can be detectable at first. Several reference-free approaches (Min-K% (Shi et al., 2023), Max-K% (Maini et al., 2024), and LOSS (Carlini et al., 2021)) can detect SFT contamination at a certain level, achieving AUROC around 73.42% across six contaminated benchmarks. The reference-based detection approach, LiRA (Mireshghallah et al., 2022), which assumes access to the training data distribution, also demonstrates superior performance with an average AUROC of 89.13% across six benchmarks.
GRPO conceals contamination. After applying GRPO to the SFT-contaminated model, we observe a consistent decrease in AUROC across all detection methods and benchmarks. We further analyze the average log probability of member and non-member samples before and after GRPO training. Fig.3 shows two key patterns: (1) GRPO lowers the entropy of generated sequences, indicating that the model becomes more confident in its generation, which is consistent with prior observations in (Cui et al., 2025); (2) the log probability distribution of members and non-members converge after GRPO. Since the gaps in log probability are the core statistical backbone of existing contamination detectors, these findings suggest that GRPO may inherently suppress contamination evidence by rendering members and non-members indistinguishable.
[…]
(a) Results on GPQA-Diamond
(b) Results on OlympiadBench
Figure 3: Log-prob distributions for members vs. non-members of SFT contaminated model before and after RL training. After further GRPO with clean samples on the SFT contaminated model, the log-prob distributions of members and non-members become increasingly similar. Since many contamination detection methods rely on separability in this space, the shrinking gap explains their degraded effectiveness. More log-prob distributions can be found in Fig. 56, and 7.
@@ 3.2 Theoretic Analysis
Training
Objectives
Clipping
Olypaid GPQA AIME25 AIME24 Minerva AMC23 Avg. $\Delta$
@@ 3.2.1 Empirical Support
To confirm empirically the prediction of our theoretical results, we evaluate the Loss detector (Carlini et al., 2021) after training with RAFT (Dong et al., 2023)/RAFT++ (Xiong et al., 2025)/GRPO. The overall results can be found in table 3. From the results, there are several observations.
@@ 4 Contamination with CoT on advanced LRMs Barely Leaves Evidence (Stage II: post-LRM)
Contamination Setup. In this setup, we simulate contamination with CoT applied to advanced LRMs at the final stage of training. We use DeepSeek-R1-Distill-Llama-8B, DeepSeek-R1-Distill-Qwen-7B (Guo et al., 2025), and checkpoints from OpenThought3 (Guha et al., 2025) as the initial models. We simulate extensive contamination with CoT by applying SFT exclusively on the member data in this section. Additional implementation details are provided in Appendix D.4.
Extensive SFT Contamination with CoT results in a huge performance inflation. As shown in Tab. 4, LRMs can substantially benefit from extensive contamination with CoT. For instance, DeepSeek-R1-Distill-Llama-8B model exhibits an average performance inflation of up to 11.76% across six benchmarks after being exposed to membership data seven times. Such inflation enables contaminated LRMs to artificially boost performance in benchmarks and have an overrated rank in the reasoning leaderboard with little extra training cost.
@@
Verbatim (Wu et al., 2025) DS Llama 48.73 50.45 41.33 61.56 59.10 40.63 50.30
DS Qwen 46.87 55.85 60.44 68.89 56.87 50.63 56.59
OpenThink 43.78 55.36 60.89 56.67 51.78 42.38 51.81
CDD (Dong et al., 2024) DS Llama 51.84 53.83 60.00 53.11 58.08 57.50 55.73
DS Qwen 51.46 48.29 50.00 53.78 54.71 41.00 49.87
OpenThink 49.98 50.23 53.31 51.24 54.52 50.44 51.62
@@
Neighbor (Mattern et al., 2023) DS Llama 49.94 39.32 53.11 43.33 49.68 60.00 49.23
DS Qwen 52.99 40.29 62.44 49.33 55.34 54.87 52.54
OpenThink 53.76 42.95 34.00 42.22 52.89 51.50 46.22
@@
LiRA (Mireshghallah et al., 2022) DS Llama 57.92 53.01 53.56 75.33 69.44 58.75 61.34
DS Qwen 46.52 43.93 50.22 58.89 59.33 54.00 52.15
OpenThink 62.35 64.77 58.44 64.44 64.81 61.62 62.74
Ref (Carlini et al., 2021) DS Llama 53.79 46.50 46.44 64.00 63.57 51.25 54.26
DS Qwen 53.30 44.37 46.89 44.22 53.09 41.75 47.27
OpenThink 57.34 49.86 37.56 50.44 59.30 69.12 53.94
@@
Zlib (Carlini et al., 2021) DS Llama 49.52 54.74 64.22 37.11 45.97 47.12 49.78
DS Qwen 46.52 57.38 64.89 36.89 43.30 42.12 48.52
OpenThink 45.65 55.37 74.22 36.89 43.51 36.62 48.71
Min–K%++ (Zhang et al., 2024) DS Llama 55.45 59.10 45.95 70.22 60.89 57.50 58.19
DS Qwen 48.92 56.83 48.44 59.33 51.83 62.62 54.66
OpenThink 51.85 58.31 66.44 55.00 49.41 41.05 53.68
Min–K% (Shi et al., 2023) DS Llama 57.86 61.68 53.33 72.67 67.12 61.87 62.42
DS Qwen 49.75 53.93 51.78 61.56 54.50 56.75 54.71
OpenThink 53.52 57.19 60.44 57.56 54.83 47.37 55.15
Max–K% (Maini et al., 2024) DS Llama 53.85 55.96 50.67 60.44 59.22 52.50 55.44
DS Qwen 49.65 50.92 40.44 73.33 54.08 56.25 54.11
OpenThink 55.12 58.29 46.22 79.33 54.20 59.38 58.76
Loss (Carlini et al., 2021) DS Llama 57.91 61.78 52.89 73.56 67.00 62.38 62.59
DS Qwen 49.77 54.09 52.00 63.78 54.76 56.75 55.19
OpenThink 53.44 57.61 61.33 56.67 55.07 48.12 55.37
@@
Figure 4: Log-prob distributions for members vs. non-members of advanced LRMs before and after SFT contamination. After extensive SFT contamination on members, the log prob of both members and non-members increases at a similar margin. More figures are in Fig. 8 and 9.
Discussion. Despite Feng et al. (2024) demonstrating that contamination detection could work in non-reasoning model scenarios, the detectors do not have access to the reasoning trajectories used in the LRM contamination scenario, so they have to rely on the generated responses from LRMs. However, LRMs typically possess strong reasoning abilities to output step-by-step long CoT and are difficult to converge on a specific sequence after contamination with long CoT. This may indicate that rather than memorizing specific reasoning trajectories, LRMs internalize the underlying knowledge and reasoning process during the contamination with CoT data, enabling generalization to distributionally similar questions (e.g., non-members). However, existing detection approaches typically assume that contaminated models assign lower loss to training sequences than to unseen ones. So, due to this generalization ability, LRMs can also achieve low loss on unseen samples that resemble the training set. This fundamentally challenges the assumption that all the detection approaches rely on, which is that benchmark contamination is more about memorizing the benchmark samples (Morris et al., 2025; Hayes et al., 2025).

### v2

@@ 1 Introduction
Accordingly, various benchmark contamination detection methods have been proposed to determine whether specific benchmarks were used during training (Yeom et al., 2018; Mattern et al., 2023; Shi et al., 2023; Dong et al., 2024; Tu et al., 2024), based on the assumption that contamination in LLMs primarily involves memorizing the benchmark data (Wu et al., 2025). These methods rely on separability in some distributions between members (i.e., seen samples during contamination) and non-members (i.e., unseen samples). However, as LLMs have started to evolve into Large Reasoning Models (LRMs) (Guo et al., 2025; Jaech et al., 2024), benchmark contamination detection faces two key challenges: (1) LRMs rely on chain-of-thought (CoT) reasoning to reach final answers, but model developers would not release their training CoT data, and contamination detectors typically only have access to question-answer pairs without the intermediate reasoning steps used during training. This absence of training sequences makes detection substantially more challenging. (2) LRMs primarily acquire reasoning abilities during two stages: SFT and RL. This potentially provides developers with opportunities to manipulate leaderboard performance by strategically contaminating benchmarks in the earlier stage (e.g., SFT), while evading detection methods through subsequent training (e.g., RL). Given these challenges, the effectiveness of existing detection methods against LRM contamination remains uncertain.
@@ 2 Related Works
Benchmark Contamination Concealment. Model developers hope to conceal contamination evidence while still having performance inflation (Dominguez-Olmedo et al., 2024). Prior work has explored evading detection through benchmark augmentation, such as rephrasing solutions with strong LLMs (Dekoninck et al., 2024; Samuel et al., 2024), but in LRM settings, most benchmarks only have question–answer pairs without step-by-step solutions, making such methods inapplicable. (Bordt et al., 2024) explores from the training dynamic perspective, showing that performance inflation due to contamination diminishes as pre-training progresses. To our knowledge, we are the first to investigate contamination concealment at the algorithmic level.
@@ 3.1 GRPO Conceals Benchmark Contamination
Contamination Inflation Mainly Comes From SFT. We evaluate multiple contamination scenarios that may happen during SFT and RL and summarize the empirical results in Tab. 1. Results show that clean SFT training yields an 11.30% improvement in pass@1 performance, while SFT contamination further inflates results by an additional 8.82% on average across six benchmarks when starting with Qwen2.5-7B-Instruct. In contrast, RL contamination, despite introducing the benchmark questions and giving rewards based on the model-generated responses, shows no significant difference compared to using a clean RL training set after short training steps.
To understand whether current contamination detection methods can still successfully detect contamination in LRMs, and whether RL training can alter the signals exploited by contamination detectors, we evaluate SFT-contaminated models before and after GRPO. Tab. 2 reveals systematic shifts in AUROC across diverse detection methods. Our analysis highlights four key observations:
SFT contamination can be detectable at first. When starting with Qwen2.5-7B-Instruct, several reference-free approaches (Min-K% (Shi et al., 2023), Max-K% (Maini et al., 2024), and LOSS (Carlini et al., 2021)) can detect SFT contamination at a certain level, achieving AUROC around 73.42% across six contaminated benchmarks. The reference-based detection approach, LiRA (Mireshghallah et al., 2022), which assumes access to the training data distribution, also demonstrates superior performance with an average AUROC of 89.13% across six benchmarks. Similar results have already been observed when starting with Llama-3.1-8B-Instruct.
Table 1: Pass@1 (%) under different contamination scenarios when the base model evolves into LRMs. The empirical results show that SFT contamination can easily inflate the benchmark performance. “/” means not used, and “Mem” denotes members. We first train the base model with SFT and then RL. The row with both “/” in the SFT/RL data columns is the result of the base model.
@@
Base model: Llama-3.1-8B-Instruct
Clean & Mem Clean & Mem 44.30 43.18 25.42 24.58 35.20 61.25 38.99
Clean & Mem Clean 44.07 48.48 27.78 25.56 37.32 66.88 41.68
Clean & Mem / 46.07 42.80 26.67 26.67 35.20 66.67 40.68
Clean Clean & Mem 44.54 40.74 25.83 23.33 29.53 61.56 37.59
Clean Clean 42.81 37.37 18.33 19.17 30.15 64.38 35.37
Clean / 40.69 39.23 16.67 18.33 27.70 56.88 33.25
/ / 15.63 29.67 0.00 4.17 19.49 19.00 14.66
Table 2: AUROC (%) of contamination detection approaches evaluated starting from an SFT-contaminated model w/o RL to subsequently trained with GRPO. Results demonstrate that after GRPO, AUROC decreases across all the benchmarks and detection approaches. $\Delta$ measures the difference with the SFT-contaminated model w/o RL. Higher AUROC, better detection performance. Each AUROC is averaged over detection scores from 8 rollouts. The base model here is Qwen2.5-7B-Instruct. More results of Llama-3.1-8B-Instruct as the base model are shown in Tab.7.
@@
GRPO conceals contamination. After applying GRPO to the SFT-contaminated model, we observe a consistent decrease in AUROC across all detection methods and benchmarks. We further analyze the average log probability of member and non-member samples before and after GRPO training, selecting Qwen2.5-7B-Instruct as the base model. Fig.3 shows two key patterns: (1) GRPO lowers the entropy of generated sequences, indicating that the model becomes more confident in its generation, which is consistent with prior observations in (Cui et al., 2025); (2) the log prob distribution of members and non-members converge after GRPO. Since the gaps in log prob are the core statistical backbone of existing contamination detectors, these findings suggest that GRPO may inherently suppress contamination evidence by rendering members and non-members indistinguishable.
@@ 3.2 Theoretic Analysis
[…]
(a) Results on GPQA-Diamond
(b) Results on OlympiadBench
Figure 3: Log-prob distributions for members vs. non-members of SFT contaminated model before and after RL training. After further GRPO with clean samples on the SFT contaminated model, the log-prob distributions of members and non-members become increasingly similar. Since many contamination detection methods rely on separability in this space, the shrinking gap explains their degraded effectiveness. More log-prob distributions can be found in Fig. 78, and 9.
@@
Training Objectives Clipping Olypaid GPQA AIME25 AIME24 Minerva AMC23 Avg. $\Delta$
@@ 3.2.1 Empirical Support
To confirm empirically the prediction of our theoretical results, we evaluate the Loss detector (Carlini et al., 2021) after training with RAFT (Dong et al., 2023)/RAFT++ (Xiong et al., 2025)/GRPO. The overall results can be found in table 3. We conduct the ablation study using Qwen2.5-7B-Instruct as the base model. From the results, there are several observations.
@@ 4 Contamination with CoT on advanced LRMs Barely Leaves Evidence (Stage II: post-LRM)
DeepSeek-R1-Distill-Qwen-14B 59.89 56.69 44.44 62.78 42.28 92.92 59.83
$\hookrightarrow$ w/ extensive SFT Contamination 66.37 62.75 64.58 77.78 46.14 97.81 69.24
@@
Verbatim (Wu et al., 2025) DS Llama-8B 48.73 50.45 41.33 61.56 59.10 40.63 50.30
DS Qwen-7B 46.87 55.85 60.44 68.89 56.87 50.63 56.59
OpenThink-7B 43.78 55.36 60.89 56.67 51.78 42.38 51.81
DS Qwen-14B 48.51 50.73 52.38 61.11 55.18 53.79 53.62
CDD (Dong et al., 2024) DS Llama-8B 51.84 53.83 60.00 53.11 58.08 57.50 55.73
DS Qwen-7B 51.46 48.29 50.00 53.78 54.71 41.00 49.87
OpenThink-7B 49.98 50.23 53.31 51.24 54.52 50.44 51.62
DS Qwen-14B 55.82 45.50 43.11 46.67 56.13 56.45 50.61
@@
Neighbor (Mattern et al., 2023) DS Llama-8B 49.94 39.32 53.11 43.33 49.68 60.00 49.23
DS Qwen-7B 52.99 40.29 62.44 49.33 55.34 54.87 52.54
OpenThink-7B 53.76 42.95 34.00 42.22 52.89 51.50 46.22
DS Qwen-14B 53.20 42.23 50.89 44.00 53.46 57.38 50.19
@@
LiRA (Mireshghallah et al., 2022) DS Llama-8B 57.92 53.01 53.56 75.33 69.44 58.75 61.34
DS Qwen-7B 46.52 43.93 50.22 58.89 59.33 54.00 52.15
OpenThink-7B 62.35 64.77 58.44 64.44 64.81 61.62 62.74
DS Qwen-14B 59.93 55.23 75.56 66.00 66.55 70.00 65.55
Ref (Carlini et al., 2021) DS Llama-8B 53.79 46.50 46.44 64.00 63.57 51.25 54.26
DS Qwen-7B 53.30 44.37 46.89 44.22 53.09 41.75 47.27
OpenThink-7B 57.34 49.86 37.56 50.44 59.30 69.12 53.94
DS Qwen-14B 55.75 47.55 52.67 30.89 55.51 53.75 49.35
@@
Zlib (Carlini et al., 2021) DS Llama-8B 49.52 54.74 64.22 37.11 45.97 47.12 49.78
DS Qwen-7B 46.52 57.38 64.89 36.89 43.30 42.12 48.52
OpenThink-7B 45.65 55.37 74.22 36.89 43.51 36.62 48.71
DS Qwen-14B 48.12 56.71 70.44 43.56 45.92 51.50 52.71
Min–K%++ (Zhang et al., 2024) DS Llama-8B 55.45 59.10 45.95 70.22 60.89 57.50 58.19
DS Qwen-7B 48.92 56.83 48.44 59.33 51.83 62.62 54.66
OpenThink-7B 51.85 58.31 66.44 55.00 49.41 41.05 53.68
DS Qwen-14B 52.44 56.72 48.44 76.44 57.39 59.62 58.51
Min–K% (Shi et al., 2023) DS Llama-8B 57.86 61.68 53.33 72.67 67.12 61.87 62.42
DS Qwen-7B 49.75 53.93 51.78 61.56 54.50 56.75 54.71
OpenThink-7B 53.52 57.19 60.44 57.56 54.83 47.37 55.15
DS Qwen-14B 52.77 58.08 52.44 77.33 59.43 59.62 59.95
Max–K% (Maini et al., 2024) DS Llama-8B 53.85 55.96 50.67 60.44 59.22 52.50 55.44
DS Qwen-7B 49.65 50.92 40.44 73.33 54.08 56.25 54.11
OpenThink-7B 55.12 58.29 46.22 79.33 54.20 59.38 58.76
DS Qwen-14B 50.43 53.89 50.00 50.00 51.08 52.50 51.32
Loss (Carlini et al., 2021) DS Llama-8B 57.91 61.78 52.89 73.56 67.00 62.38 62.59
DS Qwen-7B 49.77 54.09 52.00 63.78 54.76 56.75 55.19
OpenThink-7B 53.44 57.61 61.33 56.67 55.07 48.12 55.37
DS Qwen-14B 52.81 58.39 52.89 77.56 59.37 60.37 60.23
Contamination Setup. In this setup, we simulate contamination with CoT applied to advanced LRMs at the final stage of training. We use DeepSeek-R1-Distill-Llama-8B, DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Qwen-14B (Guo et al., 2025), and checkpoints from OpenThought3 (Guha et al., 2025) as the initial models. We simulate extensive contamination with CoT by applying SFT exclusively on the member data in this section. Additional implementation details are provided in Appendix D.4.
Extensive SFT Contamination with CoT results in a huge performance inflation. As shown in Tab. 4, LRMs can substantially benefit from extensive contamination with CoT. Such inflation enables contaminated LRMs to artificially boost performance in benchmarks and have an overrated rank in the reasoning leaderboard with little extra training cost.
@@
Figure 4: Log-prob distributions for members vs. non-members of advanced LRMs before and after SFT contamination. After extensive SFT contamination on members, the log prob of both members and non-members increases at a similar margin. More figures are in Fig. 13 and 14.
Discussion. Despite Feng et al. (2024) demonstrating that contamination detection could work in non-reasoning model scenarios, the detectors do not have access to the reasoning trajectories used in the LRM contamination scenario, so they have to rely on the generated responses from LRMs. However, LRMs typically possess strong reasoning abilities to output step-by-step long CoT and are difficult to converge on a specific sequence after contamination with long CoT. This may indicate that rather than memorizing specific reasoning trajectories, LRMs internalize the underlying knowledge and reasoning process during the contamination with CoT data, enabling generalization to distributionally similar questions (e.g., non-members). While most detection methods rely on the assumption that contaminated models would achieve lower loss on training sequences (Carlini et al., 2021) or generate less diverse responses for seen questions (Dong et al., 2024) than for unseen ones. Accordingly, these methods rely on a gap in certain metrics (e.g., log-probability, Levenshtein distance, etc.) between trained and unseen samples to determine contamination. Nevertheless, these LRMs could also have lower loss when responding to those unseen samples that share similar distributions to the training set, benefiting from their long CoT ability, as shown in Fig. 4. This confounding factor (i.e, generalization) is not accounted for by existing detection approaches, challenging the assumption that benchmark data contamination is more about memorization (Wu et al., 2025; Morris et al., 2025; Hayes et al., 2025).
@@ 5 Conclusion
Acknowledgements
The authors thank Rui Yang and Yifan Sun for their helpful feedback. Huan Zhang is supported in part by the AI2050 program at Schmidt Sciences (AI2050 Early Career Fellowship).
Reproducible claim
Our code is available at https://github.com/ASTRAL-Group/LRM_Conta_Detection_Arena.git. Detailed implementation of detection approaches, SFT training, and RL training can be found in appendix D.2 and  D.4. We also provide the proof of our theory in appendix C.
Ethics statement
We find a new vulnerability of LRM evaluations: contamination introduced at either stage can be concealed. Other than this, we do not have more ethics concerns.
Appendix contents